OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A. 2 B. - 4/5 C. 5/4 D. -2 I think C

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (alexandervonhumboldt2):

use slope formula \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

substitute your values and solve

OpenStudy (alexandervonhumboldt2):

(x,y) (-5-1)/(1-4)=solve

pooja195 (pooja195):

\[\LARGE \frac{1-(-5)}{4-1}=slope\]

OpenStudy (anonymous):

-2 ?

OpenStudy (alexandervonhumboldt2):

(-6)/(-3)=6/3=???

OpenStudy (alexandervonhumboldt2):

no not -2.

pooja195 (pooja195):

@AlexandervonHumboldt2 chill let the asker reply

OpenStudy (alexandervonhumboldt2):

what do you mean?

OpenStudy (anonymous):

ooo so it would just be 2 ? , I understand what I did wrong

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

Okay Ty

OpenStudy (alexandervonhumboldt2):

@pooja195 please don't make useless remarks to me. I'm 21 and I know what to do better than you. you know rules of os not better than me.

pooja195 (pooja195):

@AlexandervonHumboldt2 age is just a number. I asked you to be patient because the asker was still trying to reply and you were trying to rush things we help people out for them to understand not rush through so that they remain confused. I apologize if you though my comment was a useless remark.

OpenStudy (alexandervonhumboldt2):

you your self didn't waited for her reply and posted a new comment

OpenStudy (alexandervonhumboldt2):

if you are `A` it doesn't mean you can command me. i was patient. everyone has different point of views. i will write a complain on you if you will ever tell me such things. mods didn't remarked me, so I'm patient.

pooja195 (pooja195):

Go ahead write a complain on me i dont really care i'm resigning as ambi anways

OpenStudy (alexandervonhumboldt2):

Mods didn't told me that i'm not patient. Thus I'm patient and askers also didn't complained that i'm not patient. You are still trying to show that i'm a baby and you are supervisor for me.